Fuzzy Button Clothing Companys income statement reports data for its first year of operation. The firms CEO would like sales to increase by 25% next year 1, Fuzzy Button is able to achieve this level of increased sales, but its interest costs increase from 10% to 15% of earnings before interest and taxes (EBIT) 2. The companys operating costs (excluding depreciation and amortization) remain at 60% of net sales, and its depreciation and amortization expenses remain constant from year to year 3. The companys tax rate remains constant at 40% of its pre-tax income or earnings before taxes (EBT) 4. In Year 2, Fuzzy Button expects to pay $150,000 and $1,466,250 of preferred and common stock dividends, respectively Complete the Year 2 income statement data for Fuzzy Button, then answer the questions that follow. Be sure to round each dollar value to the nearest whole dollar Fuzzy Button Clothing Company Income Statement for Year Ending December 31 Year 2 Year 1 (Forecasted) $25,000,000 15,000,000 1,000,000 $9,000,000 900,000 8,100,000 Net sales Less: Operating costs, except depreciation and amortization Less: Depreciation and amortization expenses 1,000,000 Operating income (or EBIT) Less: Interest expense Pre-tax income (or EBT)
Answer #1
Income Statement
Net Sales 31250000
Less:Operating costs 18750000
Depreciation 1000000
EBIT 11500000
Less: Interest 1725000
EBT 9775000
Less: Taxes 3910000
Earnings after tax 5865000
Less: Preferred stock dividend 150000
Earnings available to common stockholders 5715000
Less: Common stock dividends 1466250
Contribution to retained earnings 4248750

1:

Dividend per share = Total preference dividend/ Number of shares

= 150,000/10000

=$15

2: EPS in year 1 = Common stock dividend/ 200000 shares

= 1215000/200000

=$6.075

EPS in year 2 = 1466250/200,000 = $7.33

3: EBITDA in year 1 = EBIT+ depreciation

= 9000,000+ 1000,000

=$10,000,000

Year 2: EBITDA =11500000+ 1000000

= 12500,000

4: wrong; not all

Addition to retained earnings are not equal to cash flows because there are some non cash items such as depreciation also.